A superhero is running toward a skyscraper to stop a villain from getting away. As he is moving, he looks up at the top of the building and his line of sight makes a 33.0° angle with the ground. After having moved 350 m, he looks at the top of the building again, and notices that his line of sight makes a 52.0° angle with the horizontal. How tall is the skyscraper?A superhero is running toward a skyscraper to stop
Answer #1

\mathrm{\theta_{1}=33^{\circ}}\rightarrow \tan\theta_{1}=\tan(33^{\circ})=0.649

\mathrm{\theta_{2}=52^{\circ}}\rightarrow \tan\theta_{2}=\tan(52^{\circ})=1.279

\mathrm{X_{1}=350m+X_{2}}

first equation

\tan\theta_{1}=\mathrm{\frac{h}{X_{1}}=\frac{h}{350+X_{2}}}\rightarrow   \mathrm{h=(350+X_{2})}\tan\theta_{1}

second equation

\tan\theta_{2}=\mathrm{\frac{h}{X_{2}}}\rightarrow   \mathrm{h=X_{2}}\tan\theta_{2}

then,

\mathrm{(350+X_{2})}\tan\theta_{1}=\mathrm{X_{2}}\tan\theta_{2}

\mathrm{X_{2}=350m}\left(\frac{\tan\theta_{1}}{\tan\theta_{2}-\tan\theta_{1}}\right)=\mathrm{360.5m}

therefore,

\mathrm{h=X_{2}}\tan\theta_{2}=\mathrm{(360.5m)(1.279)=461.4m}